# Floods do not sink prices, historical memory does: How flood risk impacts the Italian housing market

### v2

@@ 4.3 Differential effects across socio-economic groups
0.003
Observations 552728 552728
R2 0.287 0.326
Spatial fix. effects OMI zone OMI zone
N. clusters 20297 20297
@@
Models 1-2 do not consider the price as an independent variable. On average across the whole Italian territory (Model 1), people buying a home in a flood-risk area have an income 0.9% lower than people buying properties in non-risk areas. Model 2 highlights that buyers’ income differences are larger than the national average (and significant) for high awareness transactions while this effect becomes smaller for medium and low awareness transactions. Once the price correction is applied (Models 3-4), the Italian average risk coefficient (Model 3) decreases to -0.5%. This result suggests that people with lower income mostly buy homes at risk because they are cheaper, but other factors related to the flood risk contribute to explaining the income differences.
@@ 5 Conclusion
This is the first work that systematically analyzes the impact of flood risk on home prices throughout Italy. Our study reveals three key findings. First, unlike much of the existing literature, we find that even the largest recent flood in Italy–the 2023 Emilia-Romagna flood–did not lead to additional price declines for at-risk homes that were not directly hit. Second, while homes in flood-prone areas generally sell at a small discount of about 1%, this penalty is concentrated in regions with high awareness conditions (i.e., frequently flooded), where at-risk homes sell at up to 4% less than safe homes. Third, experienced buyers obtain a discount when purchasing at-risk properties, whereas only high-income young buyers do so. Low- and middle-income young buyers purchase at-risk homes at premium. Moreover, lower-income buyers are more likely to settle in flood-prone areas, particularly in the regions with high awareness, suggesting a socioeconomic sorting effect.
To make sense of these findings, we hypothesize that it is the repeated exposure to floods, rather than isolated events, what ultimately leads to the capitalization of flood risk in home prices. We argue that, with no institutional requirements for flood risk disclosure or insurance that encourages the sharing of information regarding flood risk, households rely on the historical memory of major floods to identify the most at-risk areas. This explains why price penalties emerge only for transactions under high awareness conditions and why experienced buyers–but not younger ones–obtain price discounts for at-risk properties. Finally, we find that lower-income buyers are more likely to settle in flood-prone areas, highlighting potential socioeconomic disparities in climate risk exposure, especially for the next years.
@@ B.1 Specific flood events
Figure A4 reports the robustness analysis for the 2023 Emilia-Romagna flood (Equation 3). In this case, we tested only OMI microzone and Census tract as spatial fixed effects, resulting in 24 parameter configurations. [^47] We omitted the municipality fixed effects because they are already considered in the definition of the “hit” municipalities and so including the fixed effects would lead to collinearity. The results for the selected configuration agree with those for the other specifications, supporting the results reported in the main text.
@@
Code Location Date
# homes
in aff. mun.
# homes at
risk
# hit
homes
@@ B.2 Hedonic regressions
Table A6: Coefficients for all the variables in the baseline hedonic regression (first column of Table 2).
@@
Half life
time ($\tau$)
7 10 17 7 10 17 7 10 17

### v3

@@ 4.3 Low-income buyers are more likely to move to at-risk areas
Observations 552728
R2 0.287
Spatial fix. effects OMI zone
N. clusters 20297
@@
Figure 6: Flood risk effect on buyers’ income disaggregated across awareness and mortgage applicant age, Equation (8).
@@ 5 Conclusion
This is the first work that systematically analyzes the impact of flood risk on home prices throughout Italy. Our study reveals three key findings. First, unlike much of the existing literature, we find that even the largest recent flood in Italy–the 2023 Emilia-Romagna flood–did not lead to additional price declines for at-risk homes that were not directly hit. Second, while homes in flood-prone areas generally sell at a small discount of about 1%, this penalty is concentrated in regions with high awareness conditions (i.e., frequently flooded), where at-risk homes sell at up to 4% less than safe homes. Third, only high-income buyers obtain a discount when purchasing at-risk properties, while lower-income buyers are more likely to settle in flood-prone areas, suggesting a socioeconomic sorting effect.
To make sense of these findings, we hypothesize that it is the repeated exposure to floods, rather than isolated events, what ultimately leads to the capitalization of flood risk in home prices. We argue that, with no institutional requirements for flood risk disclosure or insurance that encourages the sharing of information regarding flood risk, households rely on the historical memory of major floods to identify the most at-risk areas. Finally, we find that lower-income buyers are more likely to settle in flood-prone areas, highlighting potential socioeconomic disparities in climate risk exposure, especially for the next years.
@@ B.1 Specific flood events
Figure A4 reports the robustness analysis for the 2023 Emilia-Romagna flood (Equation 3). In this case, we tested only OMI microzone and Census tract as spatial fixed effects, resulting in 24 parameter configurations. [^47] We omitted the municipality fixed effects because they are already considered in the definition of the “hit” municipalities and so including the fixed effects would lead to colinearity. The results for the selected configuration agree with those for the other specifications, supporting the results reported in the main text.
@@
Code Location Date # homes in aff. mun. # homes at risk # hit homes
@@ B.2 Hedonic regressions
Table A6: Coefficients for all the variables in the baseline hedonic regression (third column of Table 2).
@@
Half time memory ($\tau$) 7 10 17 7 10 17 7 10 17